We are evaluating a litter removal policy with a marginal benefit function of MB=50-Q and a...

  1. We are evaluating a litter removal policy with a marginal benefit function of MB=50-Q and a marginal cost function of MC=10+Q, where Q is the number of miles of highway cleaned.
    1. Find the optimal level of clean up in this market.
    2. Find the net benefits of cleaning up 10 miles. Does that pass a cost benefit test?
    3. Find the net benefits of cleaning up 25 miles. Does that pass a cost benefit test?
Answer #1

a)

Marginal Net benefit=MNB=MB-MC=50-Q-(10+Q)=40-2Q

To find the optimal level of clean up, put MNB=0

40-2Q=0

Q=20 miles (optimal level of clean up)

b)

Total benefit (TB) of cleaning up to 10 miles is given by

Total cost (TC) of cleaning up to 10 miles is given by

Net benefit=TB-TC=450-150=300 (units)

B-C Ratio=TB/TC=450/150=3

B-C Ratio is higher than 1, it means that it passes cost benefit test.

c)

Total benefit (TB) of cleaning up to 25 miles is given by

Total cost (TC) of cleaning up to 25 miles is given by

Net benefit=TB-TC=937.50-562.50=375 (units)

B-C Ratio=TB/TC=937.50/562.50=1.67

B-C Ratio is higher than 1, it means that it passes cost benefit test.